show all clear steps.
Find Laplace Transform of the following function showing all the stepso fct) = (t + et cash (2t)) it. sinklets of
Answer #1

f(t) = (t + est cosh(2+)]{+ t sinh at + 7+ $ l (flt)) = Il + tedt coshat) + llt sinh2+) + L17) Now solving llt + e 3t cosh at2 7h Sulostituting value of coshat i 0, TH+ est cosh at) elt te* (* te>))* --**4 + 1** = (st 11 te)* - (at + D? + 2*8* + dKnow S 4 Now lt sinh2t), We know that L (sin b 2 t) : Llit sinbat) CD lau) Definuje 20 -25 L (5²-4) ² S -4 [ 1 [sinhat) - Ll+From © © ¢ 217 Lt)) = + 1 + 1 + 1 + 1 a +45 + Ś2 - 41S+8) (sty)2 + + 21st4) 45 (5²-4) ² + 1 s-107 1